OpenStudy (anonymous):

WRITE THE equation in slope intercept form of the line that is PERPENDICULAR to the graph of each equation and passes through the given point. Y=2X-3;(-5,3)

OpenStudy (anonymous):

ANYONE?!??!

OpenStudy (anonymous):

Slope intercept form is: y= mx + b Since it's perpendicular, the slope is opposite. For example: perpendicular to y = 2x - 1 through (3,4) y - 4 = -1/2 (x - 3)

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

okay. yes i do..i just get confused how to set up the equation.

OpenStudy (anonymous):

Oh alright. That's fine. Just remember since your x in (-5,3) is a negative it would be an addition sign instead of a subtraction sign \[y-y _{1} = m (x + x _{1})\]

OpenStudy (anonymous):

yes. i got that. then?

OpenStudy (anonymous):

Just plug things in \[y - 3 = -\frac{ 1 }{ 2 } ( x + 5 )\] Then distribute -1/2 to (x + 5) and tell me what you get.

OpenStudy (anonymous):

okay i have my finally answer... i got Y=-1/2+0.5? thats what i got for my finally answer? right or wrong???

OpenStudy (anonymous):

That's correct (:

OpenStudy (anonymous):

THANKKKKK YOUUUU!

OpenStudy (anonymous):

You're welcome.